OpenStudy (anonymous):

LOGARITHMS help please D; 1. e^x-3 = 10^4-x 2. 5^2x+1 = 2^4x-3

OpenStudy (anonymous):

e^(x-3) = 10^(4-x) ln(e^(x-3)) = ln(10^(4-x)) // ln(e) = 1 (x-3) = (4-x)*ln(10) // ln(10) = 2.3 x-3 = ln(10)*4-x*ln(10) x-3 = (2.3*4)-2.3x x-3 = 9.2-2.3x 3.3x = 12.2 x = 3.696969696

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

your welcome